OpenStudy (anonymous):

A triangle has side lengths of 10, 24, and 30. What type of triangle is it? a) acute b) equiangular c) right d) obtuse

OpenStudy (anonymous):

its not b and c for sure

OpenStudy (anonymous):

i was thinking d?

OpenStudy (anonymous):

is it like multiple choice where you can only have 1 answer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

oh ok according to the formula a^2 + b^2 = c^2 it would be right triangle, a^2 + b^2 > c^2 then its acute, and a^2 + b^2 < c^2 then its obtuse

OpenStudy (anonymous):

so 10^2 + 24^2 = 676 which is less than 30^2 = 900

OpenStudy (anonymous):

it should be obtuse then

OpenStudy (anonymous):

thank you :)